OpenStudy (anonymous):

write the equation in slope-intercept form. what are the slope and y-intercept? -12x+11y=-8 a. y=-12/11x+8/11 b.y=12/11; y-int 8/11 c. y=12/11x-8/11 d.y=12/11x+8/11

pooja195 (pooja195):

First we need to put the problem in slope intercept form -12x+11y=-8 Isolate y Start by adding 12x to both sides What are you left with?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

\[\huge~\rm~\bf~-12x+12x+11y=-8+12x\]

pooja195 (pooja195):

So you would have 11y=12x-8 now divide everything by 11

OpenStudy (anonymous):

y=12x-8/11

pooja195 (pooja195):

y=12/11x-8/11 y=mx+b m=slope b=y intercept

OpenStudy (anonymous):

thank you for helping me

pooja195 (pooja195):

yw ^_^